First National Bank of Bannerville has posted interest revenues of $75 million and interest costs from all of its borrowings of $45 million.

(1) If this bank possesses $800 million in total earning assets, what is First National’s net interest margin?

(2) Suppose the bank’s interest revenues and interest costs double, while its earning assets increase by 40 percent. What will happen to its net interest margin?
Answer #1

(1) Interest Revenue = $ 75 million, Interest Costs = $ 45 million

Net Interest = 75 - 45 = $ 30 million

Earnings Asset = $ 800 million

Net Interest Margin = (30/800) x 100 = 3.75 %

(2) If both interest revenue and interest costs double, the net interest will also double to become (30 x 2) = $ 60 million

Increment in Earnings Asset = 40 %

Total Earnings Asset = 1.4 x 800 = $ 1120 million

Net Interest Margin = (60/1120) x 100 = 5.35714 % ~ 5.36 %